OpenStudy (anonymous):

2x + 5y = -9 3x - 4y = -2

pooja195 (pooja195):

question ?

TheSmartOne (thesmartone):

Noted.

OpenStudy (anonymous):

Type the ordered pair that is the solution to these equations.

pooja195 (pooja195):

-3(2x + 5y = -9) 2(3x - 4y = -2) opposites cancel therefore if we multiply by -3 and 2 we will get opposite numbers a - and a +

pooja195 (pooja195):

-3(2x + 5y = -9) 2(3x - 4y = -2) distribute

OpenStudy (anonymous):

what is the answer

pooja195 (pooja195):

im not giving you the answer. Do the work.

TheSmartOne (thesmartone):

Welcome to OpenStudy! @haley.phillips We guide you to the answer; We do not give you the answer. So do what Pooja told you to do, and you'll be moving closer to your answer. And you'll understand how to do such questions :)

OpenStudy (anonymous):

x+4y=-30 2x-5y=72 --- put the system of linear equations into standard form --- 3x+4y=-30 2x-5y=72 --- copy and paste the above standard form linear equations in to this solver: https://sooeet.com/math/system-of-linear-equations-solver.php --- solution: 3x = 6